Az első valószínűségszámítás dolgozat eredménye

1.  2.  3.  4.  5.  6.

diákok:

Angyal Dóra  ±−  +  −  −  +  −
Bakó Ágnes  −  ±  −  −  −  ±−
Bathó Gábor  −  −  −  −  ±  −
Bócsi Judit  −  −  −  ±+  ±  −
Gyöngyösi Levente  −  −  −  −  ±+  −
Kovács Viktor  ±−  +  −  −  −  ±−
Ködöböcz József  −  −  −  −  ±+  −
Nagy Beáta  −  +  −  +  +  −
Nagy Eszter  −  −  +  −  +  −
Papp Ágnes  −  −  −  −  −  −
Pósa Csenge  −  −  −  −  −  −
Szeifert Diana  −  +  −  −  −  −
Tóth Veronika  −  +  −  ±  −  −